# appendices

### v1

@@ 5.4.2 ICL-Enhanced Relation Prediction
5.5 RQ4: What is the efficiency of CTINexus?
In this RQ, we assess the average token and time costs of three modules within CTINexus, using GPT-3.5 and GPT-4 as backbone models. The results, shown in Table X, indicate that using GPT-4 as the backbone results in token costs 20-30 times higher than those of GPT-3.5. Additionally, the time cost of using GPT-4 is approximately twice as high compared to GPT-3.5 for each module and the overall pipeline. The ICL-enhanced relation prediction module is the most computationally expensive, requiring multiple inferences for each input CTI. In contrast, the cybersecurity triplet extraction and hierarchical entity alignment modules have similar token costs, approximately half that of the long-distance relation prediction module, as they adhere to the “one input, one inference” principle, making them more economical. Specifically, for the hierarchical entity alignment module, the token and time costs are mainly attributed to the coarse-grained entity grouping module. The fine-grained entity merging module, which uses the text-embedding-3-large model, incurs minimal costs (\$0.13 per 1M tokens), resulting in the entire experiment costing less than \$0.30.
@@ 6 Discussion
In CTINexus, the demonstrations must be carefully chosen and of high quality, with correct answers and the required prompt format. This ensures that CTINexus can fully utilize the ICL capability to infer the correct answers from the provided examples. According to Zhao et al. [85], CTINexus’s performance degrades significantly if the demonstration set contains incorrect or misformatted samples. Additionally, although CTINexus can operate in a data-constrained manner, it still requires a certain amount of labeled data, with a recommended minimum of 100 samples. Data imbalance within the demonstration set also affects CTINexus’s performance, as an imbalanced label distribution leads to less diverse retrieved examples, increasing the likelihood of biased content generation and reducing overall effectiveness.
@@
Large Language Models (LLMs) can generate hallucinations, which are plausible yet factually inaccurate outputs [42, 76]. For instance, CTINexus with GPT-3.5 extracted the incorrect triplet $\langle$“July 2022”, “threat actors behind FARGO attacks were hijacking”, “vulnerable Microsoft SQL servers”$\rangle$ instead of $\langle$“vulnerable Microsoft SQL servers”, “are hijacked by”, “July 2022$\rangle$, leading to a complete misplacement of the subject and object and an incoherent relation. This issue is more prevalent in smaller models like GPT-3.5, LLaMA3-70B, and QWen2.5-72B. While potential solutions include fine-tuning hallucination detection classifiers or using stronger LLMs for verification, we leave these challenges for future work. Our current focus is on CSKG construction under data scarcity, where GPT-4 has demonstrated reliable performance.
@@
Various applications can be potentially empowered by CTINexus. For example, the extracted CTI knowledge can be converted into open formats like STIX [49] (e.g., also via LLMs), and exchanged in platforms like AlienVault OTX [2], and integrated into intrusion detection systems [52, 31]. A question-answering system can be developed upon the constructed CSKG using LLM’s retrieval-augmented generation [73], to provide grounded answers to threat-related questions. Cyber threat hunting [83] can also potentially be enhanced. For example, the effort required for manually constructing threat hunting queries can be reduced by using LLMs to synthesize or suggest next steps based on the constructed CSKG and partial user input. We leave the exploration of these applications for future work.
@@ 7 Related Work
Recent works have explored applying LLMs to cybersecurity challenges. PentestGPT [35] investigates LLM capabilities in penetration testing, revealing that while LLMs can handle fundamental tasks and use testing tools competently, they struggle with context loss and attention issues. TitanFuzz [36] introduces an innovative approach for fuzzing deep-learning libraries using LLMs. It employs a generative LLM for high-quality seed programs and an infilling LLM for mutations, significantly improving API and code coverage, and detects numerous previously unknown bugs. Recent studies have also explored the use of LLMs in tasks such as vulnerability detection [38, 59], patch generation [50], malware detection [29, 70], botnet traffic analysis [53, 40], and phishing and scam detection [46, 51]. Unlike these works, CTINexus leverages the ICL paradigm of LLMs for comprehensive CTI knowledge extraction and CSKG construction.
@@ 8 Conclusion and Future Work
We proposed CTINexus, a new framework leveraging ICL of LLMs for efficient and adaptive CTI extraction and CSKG construction. Unlike existing methods, CTINexus requires minimal data and parameter tuning and can adapt to various ontologies with minimal data annotation. Extensive evaluations demonstrated CTINexus’s effectiveness in extracting comprehensive knowledge, highlighting its potential to transform CTI analysis into a data-efficient and adaptable paradigm.
Future directions include integrating CTINexus with downstream applications, such as intrusion detection systems, penetration testing tools, or cybersecurity question-answering systems, to enable timely knowledge updates and ensure factual accuracy in alignment with the evolving threat landscape.
Appendix A Data Availability
Upon acceptance of this paper, all code and datasets necessary to reproduce the results will be made publicly available under an open-source license. The code and data will be shared via a publicly accessible repository to ensure transparency and facilitate further research in the field.

### v2

@@ 5.4.2 Long-Distance Relation Prediction
5.5 RQ4: How well does CTINexus perform in end-to-end CSKG construction?
In contrast to RQ2 and RQ3, which evaluate the effectiveness of each CTINexus’s component using ground truth inputs, this research question aims to assess the end-to-end performance of CTINexus, where each component operates on the output of the preceding one. We evaluated the final CSKGs constructed by CTINexus in the end-to-end setting by comparing them to the ground truth CSKGs. The evaluation used the triplet-level metrics described in Section 5.1, where a triplet was considered correct if its subject, relation, and object semantically match those of a gold-standard triplet. We evaluated CTINexus with its optimal configuration on ten sampled CTI reports from our dataset, and achieved an end-to-end F1-Score of 87.80%, a precision of 81.82%, and a recall of 94.74%. These findings confirm that CTINexus exhibits minimal error propagation across phases, ruling out the snowball effect and ensuring robust utility for downstream applications.
5.6 RQ5: How well does CTINexus adapt to different CSKG ontologies?
To evaluate the adaptability of CTINexus to various CSKG ontologies, we conducted an experiment using the STIX ontology, a widely recognized threat intelligence-sharing standard [22] supported by numerous vendors. STIX categorizes entities into STIX Domain Objects (SDOs) and STIX Relationship Objects (SROs) to systematically capture entities and their interrelationships within threat intelligence. For our evaluation, we selected 13 SDOs: Campaign, Grouping, Identity, Indicator, Infrastructure, Intrusion Set, Location, Attack Pattern, Malware, Threat Actor, Tool, Vulnerability, and Report. Five SDOs, Course of Action, Note, Opinion, Malware Analysis, and Observed Data, were omitted because they are usually paragraph-length text blocks that bundle multiple entities, whereas CSKG nodes should represent discrete, meta-level concepts. We performed an experiment on a small-scale corpus of ten CTI reports (mean length $\approx 1,318$ tokens) from STIXnet [23]. Manual annotation under the STIX ontology yielded 128 gold-standard triplets. On this set, CTINexus achieved 89.7 % precision, 81.9 % recall, and an 85.6 % F1-score—only a few points below its MALOnt performance. This small drop confirms CTINexus’s resilience to ontology shifts: the underlying LLM weights remain frozen, and the target ontology is simply injected into the prompt (and mirrored in a handful of ontology-consistent demonstration examples), enabling the model to adapt on-the-fly to new ontologies.
5.7 RQ6: What is the efficiency of CTINexus?
We assessed the average token and time costs of three modules within CTINexus, using GPT-3.5 and GPT-4 as backbone models. As shown in Table X, the average cost per CTI report is \$0.1485 with GPT-4 and only \$0.0069 with GPT-3.5. The maximum per-report cost reaches \$0.1778 (GPT-4) and \$0.0086 (GPT-3.5), while the minimum costs are \$0.0423 and \$0.0021, respectively. The results indicate that using GPT-4 as the backbone results in token costs 20-30 times higher than those of GPT-3.5. Additionally, the time cost of using GPT-4 is approximately twice as high compared to GPT-3.5 for each module and the overall pipeline. The ICL-enhanced relation prediction module is the most computationally expensive, requiring multiple inferences for each input CTI. In contrast, the cybersecurity triplet extraction and hierarchical entity alignment modules have similar token costs, approximately half that of the long-distance relation prediction module, as they adhere to the “one input, one inference” principle, making them more economical. Specifically, for the hierarchical entity alignment module, the token and time costs are mainly attributed to the coarse-grained entity grouping module. The fine-grained entity merging module, which uses the text-embedding-3-large model, incurs minimal costs (\$0.13 per 1M tokens), resulting in the entire experiment costing less than \$0.30.
@@ 6 Discussion
In CTINexus, the demonstration examples must be carefully chosen and of high quality, with correct answers and the required prompt format. This ensures that CTINexus can fully utilize the ICL capability to infer the correct answers from the provided examples. CTINexus’s performance can degrade if the demonstration set contains incorrect or misformatted samples. Additionally, although CTINexus can operate in a data-constrained manner, it still requires a certain amount of labeled data, with a recommended minimum of 100 samples. Data imbalance within the demonstration set also affects CTINexus’s performance, as an imbalanced label distribution leads to less diverse retrieved examples, increasing the likelihood of biased content generation and reducing overall effectiveness.
@@
LLMs can generate hallucinations, which are plausible yet factually inaccurate outputs [47]. For instance, CTINexus with GPT-3.5 extracted the incorrect triplet $\langle$“July 2022”, “threat actors behind FARGO attacks were hijacking”, “vulnerable Microsoft SQL servers”$\rangle$ instead of $\langle$“vulnerable Microsoft SQL servers”, “are hijacked by”, “July 2022$\rangle$, leading to a misplacement of the subject and object and an incoherent relation. This issue is more prevalent in smaller models like GPT-3.5, LLaMA3-70B, and QWen2.5-72B. While potential solutions include fine-tuning hallucination detection classifiers or using stronger LLMs for verification, we leave these challenges for future work. Our current focus is on CSKG construction under data scarcity, where GPT-4 has demonstrated reliable performance and surpasses existing state-of-the-art CTI extraction approaches by a large margin.
@@
CTINexus has the potential to empower various defensive applications. For example, the extracted CTI knowledge can be converted (also via LLMs) into open formats like STIX [22], and exchanged in platforms like AlienVault OTX [2], and ingested by intrusion prevention systems [68, 44, 33] and log analysis frameworks [86]. A question-answering system can be developed upon the constructed CSKG using LLM’s retrieval-augmented generation [77], to provide grounded answers to threat-related questions. Cyber threat hunting [43, 64] and investigation [45, 40] systems can also be enhanced. For example, the effort required for manually constructing investigation queries can be reduced by using LLMs to synthesize or suggest next steps based on the CSKG and partial user input. We leave the exploration of these downstream applications for future work.
@@ 7 Related Work
Recent works have explored applying LLMs to cybersecurity challenges. PentestGPT [37] investigates LLM capabilities in penetration testing, revealing that while LLMs can handle fundamental tasks and use testing tools competently, they struggle with context loss and attention issues. TitanFuzz [38] introduces an innovative approach for fuzzing deep-learning libraries using LLMs. It employs a generative LLM for high-quality seed programs and an infilling LLM for mutations, significantly improving API and code coverage, and detects numerous previously unknown bugs. Recent studies have also explored the use of LLMs in tasks such as vulnerability detection [58], patch generation [83], malware detection [89], and phishing and scam detection [55]. Unlike these works, CTINexus leverages the ICL paradigm of LLMs for comprehensive CTI knowledge extraction and CSKG construction.
@@ 8 Conclusion and Future Work
We presented CTINexus, a framework leveraging the ICL of LLMs for efficient and adaptive CTI knowledge extraction and CSKG construction. Unlike existing methods, CTINexus requires minimal data and parameter tuning and can adapt to various ontologies with minimal data annotation. Extensive evaluations demonstrated CTINexus’s effectiveness in extracting comprehensive knowledge, highlighting its potential to transform CTI analysis into a data-efficient and adaptable paradigm.
The rapid adoption of LLMs in security calls for foundation models that can ingest and refresh large-scale threat intelligence in near real time. Current solutions—continual pre-training on security corpora and retrieval-augmented generation—are computationally costly, slow to incorporate new data, and struggle to reconcile heterogeneous sources. CTINexus offers an alternative: KG-augmented generation. By continuously extracting, aligning, and integrating information into evolving cybersecurity knowledge graphs, CTINexus can supply downstream security LLMs with a structured, up-to-date memory, enabling accurate, cross-source reasoning at inference time.
Another direction for future work is the integration of visual analytics into CTINexus. Visual representations of malicious activities could aid analysts in identifying behavior patterns and relationships, offering additional context for understanding threat evolution. Such tools would enhance interpretability and assist in supporting timely cybersecurity decision-making.
Acknowledgments
We would like to thank the anonymous reviewers for their constructive feedback. This work is supported in part by the Commonwealth Cyber Initiative (CCI). Any opinions, findings, and conclusions made in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.